# A lower bound on the number of edges in DP-critical graphs

### v1

@@ 4.1 General observations
so by Lemma 4.5, $j:=|E_{G}(U,\overline{U})|\leq 1$.
If $j=0$, then $U=V(G)$, every vertex of $G$ is low, and $G$ is not a GDP-tree. Therefore, $G$ is DP $h$-colorable, a contradiction. If $j=1$, then write $x$ for the unique vertex of $U$ with a neighbor outside of $U$. As $d(v)=2$, $x\neq v$. If $x\not\in\{u,w\}$, then each vertex $z\in U^{\prime}\setminus\{x\}$ is low, and $G^{\prime}[U^{\prime}\setminus\{x\}]$ is a $2$-connected graph that is not a GDP-tree. Hence, $G^{\prime}[U^{\prime}\setminus\{x\}]$ is DP $(h-d_{G}+d_{G[U^{\prime}\setminus\{x\}]})$-colorable, contradicting Observation 4.2.
Otherwise, $x\in\{u,w\}$. We claim that $G[U]$ is DP $(h-d_{G}+d_{G[U]})$-colorable. For each $z\in U$, we write $h^{\prime}(z)=h(z)-d_{G}(z)+d_{G[U]}(z)$, and we observe that $h^{\prime}(x)=k-2\geq 3$, and $h^{\prime}(z)=h(z)$ for each $z\in U\setminus\{x\}$. Consider an $h^{\prime}$-cover $(H^{\prime},L^{\prime})$ of $G[U]$. We assign a color $c\in L^{\prime}(x)$ to $x$ with no neighbor in $L(v)$; this is possible, as $|L(v)|=h^{\prime}(v)=2<|L^{\prime}(x)|$. Afterward, each vertex $z\in U\setminus\{x\}$ has at least $d_{G[U\setminus\{x\}]}(z)$ colors in $L^{\prime}(z)$ not adjacent to $c$. Furthermore, $v$ has two colors in $L(v)$ not adjacent to $c$, while $v$ has only one neighbor in $U\setminus\{x\}$. Hence, we can finish an $(H^{\prime},L^{\prime})$-coloring of $G[U]$ using degeneracy, contradicting Observation 4.2. ∎
@@
This follows from Lemmas 4.9 and 4.11. ∎
@@
Let $k\geq 5$. If $G$ is $2$-connected and $h(v)=k-1$ for each $v\in V(G)$, then $G$ is not $(k-1)$-regular.
@@
Suppose $G$ is $(k-1)$-regular. As $G$ is $h$-minimal, Theorem 1.7 implies that $G$ is either $K_{t}^{\frac{k-1}{t-1}}$ for some $22\leq t\leq k-$ 2 (satisfying $(t-1)|(k-1)$) or $C_{n}^{(k-1)/2}$ for some $n\geq 3$ (provided $k$ is odd). If $G$ is $K_{t}^{\frac{k-1}{t-1}}$, then
@@
As a function of $t$, this expression is a convex quadratic and is maximized at $t=2$. When $t=2$ and $k\geq 6$, we have $\rho_{h}(G)=2-k+2\leq-2$. Similarly, if $k\geq 6$ and $G=C_{n}^{(k-1)/2}$ (so $k$ is odd), then
@@
So suppose $k=5$. If $G=C_{n}^{2}$, then Part 2 of Theorem 2.1 holds. If $G=K_{t}^{\frac{5-1}{t-1}}$ for some $22\leq t\leq 5-$ 2, then $t\in\{2,3\}$. If $t=2$, then $G$ is isomorphic to $K_{2}^{4}$. If $t=3$, then $G$ is isomorphic to $C_{3}^{2}$. In all cases, Part 2 of Theorem 2.1 holds. ∎
4.2 A special set $S_{0}^{*}$ of vertices in $G$
An edge-block in a multigraph $G$ is an inclusion maximal connected subgraph $G^{\prime}$ of $G$ such that either $|V(G^{\prime})|=2$ or $G^{\prime}$ has no cut edges. In particular, every cut edge forms an edge-block, and each connected graph decomposes into edge-blocks.
@@ 4.2 A special set $S_{0}^{*}$ of vertices in $G$
4.3 Subgraphs induced by low vertices in $G$
Let $\mathcal{B}$ be the subgraph of $G$ induced by low vertices in $G$ and let ${\mathcal{B}}_{0}$ be the subgraph of $\mathcal{B}$ induced by the vertices of $\mathcal{B}$ in $S^{*}_{0}$. By Lemma 4.4, each block in $\mathcal{B}$ is a multiple of either a complete graph or a cycle. In particular, each such block is regular. Given a $(k-2)$-regular block $B\subseteq\mathcal{B}$ along with distinct vertices $u,u^{\prime}\in N(B)\setminus V(B)$, we write $F(B,u,u^{\prime})$ for the multigraph obtained from $G-V(B)$ by adding an edge $uu^{\prime}$.
@@ 4.3 Subgraphs induced by low vertices in $G$
If $B$ is a $(k-2)$-regular block of $\mathcal{B}$, then the multigraph $F(B,u,u^{\prime})$ is not DP $h$-colorable.
@@
As $G$ is $h$-minimal, $G-V(B)$ has an $(H,L)$-coloring $f$ that does not extend to all of $G$. For each $v\in V(B)$, we write $L_{f}(v)$ for the subset of $L(v)$ consisting of colors with no neighbor in $f$. We write $H^{\prime}=H[L_{f}(V(B))]$. As $|L_{f}(v)|\geq k-2$ for each $v\in V(B)$ and $B$ has no $(H^{\prime},L_{f})$-coloring, Lemma 3.1 implies that $(H^{\prime},L_{f})$ is one of the blowup covers described in Lemma 3.1. In particular, $H^{\prime}$ is $(k-2)$-regular. Thus, $H_{f}:=H[L(V(B))]-V(H^{\prime})$ is isomorphic to a subgraph of $\widetilde{B}$.
Now, we consider two cases. First, suppose that every component of $H[L(V(B))]$ is isomorphic to $H_{f}$. Then, as at least one vertex $v\in V(B)$ satisfies $h(v)=k-1$ (namely the neighbor of $u$ in $B$), it follows that $H[L(V(B))]$ is isomorphic to $k-1$ disjoint copies of $H_{f}$, each of which contains exactly one vertex from $L(v)$ for each $v\in V(B)$. Now, we construct an $h$-cover $(H^{\prime\prime},L)$ of $F(B,u,u^{\prime})$ as follows. We begin with $(H,L)$, and we remove $L(V(B))$ from $H$. Then, we add a perfect matching between $L(u)$ and $L(u^{\prime})$ so that $c\in L(u)$ and $c^{\prime}\in L(u^{\prime})$ are adjacent if and only if $c$ and $c^{\prime}$ are adjacent to a common component of $H[L(V(B))]$. Then, if $G-V(B)$ has an $(H^{\prime\prime},L)$-coloring $f^{\prime}$, the colors $f^{\prime}(u)$ and $f^{\prime}(u^{\prime})$ are adjacent to distinct components of $H[L(V(B))]$. Hence, there is no $H_{f}$-component in $H[L(V(B))]$ all of whose vertices are adjacent to $f^{\prime}$; therefore, by Lemma 3.1, $f^{\prime}$ can be extended to all of $G$. This contradicts the initial assumption that $G$ is $h$-minimal, completing the first case.
By Lemma 3.1, the only other case to consider is that $H[L(V(B))]$ has a single component isomorphic to $H_{f}$. In this case, $H^{\prime}$ is a $q$-blowup of either a clique or a cycle, where $q\geq 2$ is the quotient of $k-2$ and the regularity of $\widetilde{B}$. In particular, $\widetilde{B}$ has maximum degree less than $k-2$. Then, we construct an $h$-cover $(H^{\prime\prime},L)$ of $F(B,u,u^{\prime})$ as follows. We begin with $(H,L)$, and we remove $L(V(B))$ from $H$. Then, we add a single edge to $H^{\prime\prime}$ joining $f(u)$ and $f(u^{\prime})$.
Now, suppose that $F(B,u,u^{\prime})$ has an $(H^{\prime\prime},L)$-coloring $f^{\prime}$. By construction, either $f^{\prime}(u)\neq f(u)$ or $f^{\prime}(u^{\prime})\neq f(u^{\prime})$. Therefore, some vertex of $H_{f}$ has no neighbor in $f^{\prime}$. For each $v\in V(B)$, we let $L^{\prime\prime}(v)$ be the set of colors $c\in L(v)$ for which $c$ has no neighbor in $f^{\prime}$. Then, $H[L^{\prime\prime}(V(B))]$ is not $(k-2)$-regular, so by Lemma 3.1, the coloring $f^{\prime}$ extends to an $(H,L)$-coloring of $G$. This contradicts our initial assumption that $G$ is $h$-minimal.
@@
By Lemma 4.15, for each $(k-2)$-regular subgraph $B\subseteq\mathcal{B}$, and any two distinct vertices $u,u^{\prime}\in N(B)\setminus V(B)$, there is an $h$-minimal multigraph $G(B,u,u^{\prime})$ contained in $F(B,u,u^{\prime})$. As $G$ is $h$-minimal, $G(B,u,u^{\prime})$ is not a subgraph of $G$; therefore, $G(B,u,u^{\prime})$ contains $u$ and $u^{\prime}$.
@@
If $S^{*}_{0}\neq V(G)$, then the vertex $x^{*}_{0}$ is not in a ${(k-2)}$-regular block of $\mathcal{B}$.
@@
Suppose $x^{*}_{0}$ is in a ${(k-2)}$-regular block $B$ of $\mathcal{B}$. Since $B$ is $2$-connected, $y^{*}_{0}\notin V(B)$. If there exists a vertex $y_{1}^{*}\in N(V(B))\setminus(V(B)\cup\{y^{*}_{0}\})$, let $G^{\prime}=G(B,y^{*}_{0},y_{1}^{*})$. Recall that $y_{0}^{*},y_{1}^{*}\in V(G^{\prime})$. Since $y^{*}_{0}y_{1}^{*}$ is a cut edge in $G^{\prime}$, it is not part of a $K_{k}$-subgraph of $G^{\prime}$. Therefore, $\rho_{G^{\prime},h}(V(G^{\prime}))\leq-2$, and hence $\rho_{G,h}(V(G^{\prime}))\leq 2\lambda-2$. However, the multigraph $G[V(G^{\prime})]$ is disconnected. As $G[S^{*}_{0}]$ is an edge-block of $G$, $|E_{G}(V(G^{\prime})\cap S^{*}_{0},S^{*}_{0}\setminus V(G^{\prime}))|\geq 2$; therefore, by Lemma 4.5, $\rho_{G,h}(V(G^{\prime})\cap S^{*}_{0})\geq 2\lambda-3$. Hence, $\rho_{G,h}(V(G^{\prime})\setminus S^{*}_{0})\leq(2\lambda-2)-(2\lambda-3)<\lambda-1$, contradicting Observation 4.7.
Suppose now that $V(B)$ has no neighbor in $V(G)\setminus(V(B)\cup\{y^{*}_{0}\})$. Then every vertex $u\in V(B)-x_{0}^{*}$ satisfies $h(u)=k-2$. Let $b=|V(B)|$, so
@@
If $k\geq 5$, then ${\mathcal{B}}_{0}$ does not contain $K_{k-1}$ formed by $(k-1)$-vertices.
@@
Suppose there is a set $C=\{v_{1},\ldots,v_{k-1}\}\subseteq S^{*}_{0}$ consisting of low $(k-1)$-vertices such that $B:=G[C]=K_{k-1}$. For $i\in\{1,\ldots,k-1\}$, let $v^{\prime}_{i}$ be the unique neighbor of $v_{i}$ outside of $C$. Let $C^{\prime}=\{v_{1}^{\prime},\ldots,v_{k-1}^{\prime}\}$, so that $v_{i}v_{i}^{\prime}\in E(G)$ for each $i\in\{1,\dots,k-1\}$. Since $G$ has no $K_{k}$, $|C^{\prime}|\geq 2$.
For distinct $s,t\in\{1,\ldots,k-1\}$, write $G_{s,t}=G(B,v_{s}^{\prime},v^{\prime}_{t})$. We will simultaneously consider two cases: If $\rho_{h}(G_{s,t})\leq-2$ for some distinct $s,t$ or if $G_{s,t}$ is exceptional for every distinct pair $s,t$. Let $\varrho(C)=\min\{\rho_{h}(G_{s,t}):s,t\in\{1,\ldots,k-1\}\}$.
For $j\in\{1,\ldots,k-1\}$, define the function
$g(j)=\begin{cases}2\lambda+2j-5&\mbox{ if $\rho_{h}(G_{s,t})\leq-2$ for some distinct $s,t$ (Case 1)}\\ \varrho(C)+2\lambda+j-1&\mbox{ if $G_{s,t}$ is exceptional for all distinct $s,t$ (Case 2)}\\ \end{cases}$
For $j\in\{1,\dots,k-1\}$, let $\pi(j)$ be the minimum of $\rho_{G,h}(P)$ over $P\subseteq V(G)\setminus C$ with $|N(P)\cap C|\geq j$. Let $P_{j}\subseteq V(G)\setminus C$ be such that $|N(P_{j})\cap C|\geq j$ and $\rho_{G,h}(P_{j})=\pi(j)$.
We will prove that $\pi(k-1)\leq g(k-1)$. For this, let $j$ be the largest value for which $\rho_{G,h}(P_{j})\leq g(j)$. If $j=k-1$, then we are done, so assume $j<k-1$. Then $j\geq 2$ since
$\pi(2)=\rho_{G,h}(P_{2})\leq\begin{cases}-2+(2\lambda+1)=2\lambda-1=g(2).&\mbox{(Case 1)}\\ \varrho(C)+(2\lambda+1)=g(2).&\mbox{(Case 2)}\\ \end{cases}$
Assume there are distinct $i,i^{\prime}\in\{1,\ldots,k-1\}$ such that $v^{\prime}_{i}\in P_{j}$ and $v^{\prime}_{i^{\prime}}\not\in P_{j}$. Since $V(G_{i,i^{\prime}})\cup P_{j}\supseteq P_{j}\cup\{v^{\prime}_{i^{\prime}}\}$, $V(G_{i,i^{\prime}})\cup P_{j}$ has at least $j+1$ neighbors in $C$. Thus, Lemma 2.2 implies that
$\pi(j+1)\leq\rho_{G,h}(P_{j}\cup V(G_{i,i^{\prime}}))\leq g(j)+\rho_{G,h}(V(G_{i,i^{\prime}}))-\rho_{G,h}(P_{j}\cap V(G_{i,i^{\prime}})).$ (15)
@@
=$=\begin{cases}(1-\lambda)(k-1)+(2\lambda+2(k-1)-5)\leq-2.&\mbox{(Case 1)}\\ (1-\lambda)(k-1)+(\varrho(C)+2\lambda+k-1)\leq(1-\lambda)(k-1)+2\lambda+2k-2\leq-3.&\mbox{(Case 2: $k\geq 6$)}\\ (1-\lambda)(k-1)+(\varrho(C)+2\lambda+k-1)\leq(1-\lambda)(k-1)+2\lambda+k-1\leq-4.&\mbox{(Case 2: $k=5$ and $\varrho(C)\leq 0$)}\\ (1-\lambda)(k-1)+(\varrho(C)+2\lambda+k-1)=(1-\lambda)(k-1)+2\lambda+k+4=1.&\mbox{(Case 2: $k=5$ and $\varrho(C)=5$)}\end{cases}$
This contradicts Observation 4.6 in all cases except when $k=5$ and $\varrho(C)=5$. Since $11<\lambda-$ 1 when $k=5$, Observation 4.7 implies $C\cup P_{C^{\prime}}=V(G)$. In particular, $G[P_{C^{\prime}}]$ is $K_{5}^{-}$ and $G[C^{\prime}]$ is $K_{4}^{-}$ so $G_{s,t}$ is exceptional only for one pair $s,t$. Yet, this is impossible since we are in Case 2. ∎
@@
${\mathcal{B}}_{0}$ does not contain a ${(k-2)}$-regular block.

### v2

@@ 4.1. General observations
so by Lemma 4.5, $j:=|E_{G}(U^{\prime},\overline{U^{\prime}})|\leq 1$ and if $\varepsilon=1$, then $j=0$.
If $j=0$, then $U^{\prime}=V(G)$, every vertex of $G$ is low, and $G$ is not a GDP-tree. Therefore, $G$ is DP $h$-colorable, a contradiction. If $j=1$, then $\varepsilon=0$. Write $x$ for the unique vertex of $U^{\prime}$ with a neighbor outside of $U^{\prime}$. As $d(v)=2$, $x\neq v$. If $x\not\in\{u,w\}$, then each vertex $z\in U^{\prime}\setminus\{x\}$ is low, and $G^{\prime}[U^{\prime}\setminus\{x\}]$ is $2$-connected and not a GDP-tree. Hence, $G^{\prime}[U^{\prime}\setminus\{x\}]$ is DP $(h-d_{G}+d_{G[U^{\prime}\setminus\{x\}]})$-colorable, contradicting Observation 4.2.
Otherwise, $x\in\{u,w\}$. We claim that $G[U^{\prime}]$ is DP $(h-d_{G}+d_{G[U^{\prime}]})$-colorable. For each $z\in U^{\prime}$, we write $h^{\prime}(z)=h(z)-d_{G}(z)+d_{G[U^{\prime}]}(z)$, and observe that $h^{\prime}(x)=k-2\geq 3$, and $h^{\prime}(z)=h(z)$ for each $z\in U^{\prime}\setminus\{x\}$. Consider an $h^{\prime}$-cover $(H^{\prime},L^{\prime})$ of $G[U^{\prime}]$. We assign a color $c\in L^{\prime}(x)$ to $x$ with no neighbor in $L(v)$; this is possible, as $|L(v)|=h^{\prime}(v)=2<|L^{\prime}(x)|$. Afterward, each vertex $z\in U^{\prime}\setminus\{x\}$ has at least $d_{G[U^{\prime}\setminus\{x\}]}(z)$ colors in $L^{\prime}(z)$ not adjacent to $c$. Furthermore, $v$ has two colors in $L(v)$ not adjacent to $c$, while $v$ has only one neighbor in $U\setminus\{x\}$. We give a linear ordering to the vertices of $U^{\prime}\setminus\{x\}$ that ends with $v$, and observe that for each $z\in U^{\prime}\setminus\{x\}$, the number of neighbors of $z$ appearing before $z$ in the ordering is less than $|L^{\prime}(z)\setminus\{c\}|$. Therefore, by greedily coloring the vertices of $U^{\prime}\setminus\{x\}$ according to our linear ordering, we complete an $(H^{\prime},L^{\prime})$-coloring of $G[U^{\prime}]$, contradicting Observation 4.2. Hence we can finish an $(H^{\prime},L^{\prime})$-coloring of $G[U^{\prime}]$ using degeneracy, contradicting Observation 4.2. ∎
@@
This follows from Observation 4.1 and Lemmas 4.9 and 4.11. ∎
@@
Let $k\geq 5$. If $G$ is $2$-connected and h(v)=k1h(v)=k-1 for each $v\in V(G)$, then $G$ is not (k1)(k-1)-regular.
@@
Suppose $G$ is (k1)(k-1)-regular. As $G$ is $h$-minimal, Theorem 1.5 implies that $G$ is either ${\frac{k-1}{t-1}}K_{t}$ for some $2\leq t\leq k-2$ (satisfying (t1)|(k1)(t-1)|(k-1)) or ${(k-1)/2}C_{n}$ for some $n\geq 3$ (provided $k$ is odd). If $G$ is ${\frac{k-1}{t-1}}K_{t}$, then
@@
As a function of $t$, this expression is a convex quadratic function and is maximized at $t=2$. When $t=2$ and $k\geq 6$, we have $\rho_{h}(G)=2-k+2\leq-2$. Similarly, if $k\geq 6$ and $G={(k-1)/2}C_{n}$ (so $k$ is odd), then
@@
So suppose $k=5$. If $G=2C_{n}$, then Part 2 of Theorem 2.1 holds. If $G={\frac{5-1}{t-1}}K_{t}$ for some $2\leq t\leq 5-2$, then $t\in\{2,3\}$. If $t=2$, then $G$ is isomorphic to $4K_{2}$. If $t=3$, then $G$ is isomorphic to $2C_{3}$. In all cases, Part 2 of Theorem 2.1 holds. ∎
4.2. A special set $S_{0}^{*}$ of vertices in $G$
An edge-block in a multigraph $G$ is an inclusion maximal connected subgraph $G^{\prime}$ of $G$ such that either $|V(G^{\prime})|=2$ or $G^{\prime}$ has no cut edges. In particular, every cut edge forms an edge-block, and each connected multigraph decomposes into edge-blocks.
@@ 4.2. A special set $S_{0}^{*}$ of vertices in $G$
If $G[S]$ is not connected, then each vertex set $S^{\prime}$ inducing a component of $G[S]$ is a proper subset of $V(G)$, and hence $\rho_{G,h}(S^{\prime})\geq\lambda-1$ by Lemma 4.5. As $G[S]$ has at least two components, it follows that $\rho_{G,h}(S)\geq 2\lambda-2$. Thus, we assume that $G[S]$ is connected.
@@
4.3. Subgraphs induced by low vertices in $G$
Let $\Lambda$ be the subgraph of $G$ induced by low vertices in $G$ and let ${\Lambda}_{0}$ be the subgraph of $\Lambda$ induced by the vertices of $\Lambda$ in $S^{*}_{0}$. By Lemma 4.4, each block in $\Lambda$ is a multiple of either a complete graph or a cycle. In particular, each such block is regular. Given a (k2)(k-2)-regular block $B\subseteq\Lambda$ along with distinct vertices $u,u^{\prime}\in N(B)\setminus V(B)$, we write $F(B,u,u^{\prime})$ for the multigraph obtained from GV(B)G-V(B) by adding an edge $uu^{\prime}$.
@@ 4.3. Subgraphs induced by low vertices in $G$
If $B$ is a (k2)(k-2)-regular block of $\Lambda$, then the multigraph $F(B,u,u^{\prime})$ is not DP $h$-colorable.
@@
As $G$ is $h$-minimal, GV(B)G-V(B) has an $(H,L)$-coloring $\varphi$ that does not extend to all of $G$. For each $v\in V(B)$, we write $L_{\varphi}(v)$ for the subset of $L(v)$ consisting of colors with no neighbor in $\varphi$. We write $H^{\prime}=H[L_{\varphi}(V(B))]$. As $|L_{\varphi}(v)|\geq k-2$ for each $v\in V(B)$ and $B$ has no $(H^{\prime},L_{\varphi})$-coloring, Lemma 3.1 implies that $(H^{\prime},L_{\varphi})$ is one of the blowup covers described in Lemma 3.1. In particular, $H^{\prime}$ is (k2)(k-2)-regular. Thus, $H_{\varphi}:=H[L(V(B))]-V(H^{\prime})$ is isomorphic to a subgraph of $\widetilde{B}$.
Now, we consider two cases. First, suppose that every component of $H[L(V(B))]$ is isomorphic to $H_{\varphi}$. Then, as at least one vertex $v\in V(B)$ satisfies h(v)=k1h(v)=k-1 (namely the neighbor of $u$ in $B$), it follows that $H[L(V(B))]$ is isomorphic to k1k-1 disjoint copies of $H_{\varphi}$, each of which contains exactly one vertex from $L(v)$ for each $v\in V(B)$. Now, we construct an $h$-cover $(H^{\prime\prime},L)$ of $F(B,u,u^{\prime})$ as follows. We begin with $(H,L)$, and we remove $L(V(B))$ from $H$. Then, we add a perfect matching between $L(u)$ and $L(u^{\prime})$ so that $c\in L(u)$ and $c^{\prime}\in L(u^{\prime})$ are adjacent if and only if $c$ and $c^{\prime}$ are adjacent to a common component of $H[L(V(B))]$. Then, if GV(B)G-V(B) has an $(H^{\prime\prime},L)$-coloring $\varphi^{\prime}$, the colors $\varphi^{\prime}(u)$ and $\varphi^{\prime}(u^{\prime})$ are adjacent to distinct components of $H[L(V(B))]$. Hence, there is no $H_{\varphi}$-component in $H[L(V(B))]$ all of whose vertices are adjacent to $\varphi^{\prime}$; therefore, by Lemma 3.1, $\varphi^{\prime}$ can be extended to all of $G$. This contradicts the initial assumption that $G$ is $h$-minimal, completing the first case.
By Lemma 3.1, the only other case to consider is that $H[L(V(B))]$ has a single component isomorphic to $H_{\varphi}$. In this case, $H^{\prime}$ is a $q$-blowup of either a clique or a cycle, where $q\geq 2$ is the quotient of k2k-2 and the regularity of $\widetilde{B}$. In particular, $\widetilde{B}$ has maximum degree less than k2k-2. Then, we construct an $h$-cover $(H^{\prime\prime},L)$ of $F(B,u,u^{\prime})$ as follows. We begin with $(H,L)$, and we remove $L(V(B))$ from $H$. Then, we add a single edge to $H^{\prime\prime}$ joining $\varphi(u)$ and $\varphi(u^{\prime})$.
Now, suppose that $F(B,u,u^{\prime})$ has an $(H^{\prime\prime},L)$-coloring $\varphi^{\prime}$. By construction, either $\varphi^{\prime}(u)\neq\varphi(u)$ or $\varphi^{\prime}(u^{\prime})\neq\varphi(u^{\prime})$. Therefore, some vertex of $H_{\varphi}$ has no neighbor in $\varphi^{\prime}$. For each $v\in V(B)$, we let $L^{\prime\prime}(v)$ be the set of colors $c\in L(v)$ for which $c$ has no neighbor in $\varphi^{\prime}$. Then, $H[L^{\prime\prime}(V(B))]$ is not (k2)(k-2)-regular, so by Lemma 3.1, the coloring $\varphi^{\prime}$ extends to an $(H,L)$-coloring of $G$. This contradicts our initial assumption that $G$ is $h$-minimal.
@@
By Lemma 4.15, for each (k2)(k-2)-regular subgraph $B\subseteq\Lambda$, and any two distinct vertices $u,u^{\prime}\in N(B)\setminus V(B)$, there is an $h$-minimal multigraph $G(B,u,u^{\prime})$ contained in $F(B,u,u^{\prime})$. As $G$ is $h$-minimal, $G(B,u,u^{\prime})$ is not a subgraph of $G$; therefore, $G(B,u,u^{\prime})$ contains $u$ and $u^{\prime}$.
@@
If $S^{*}_{0}\neq V(G)$, then the vertex $x^{*}_{0}$ is not in a (k2){(k-2)}-regular block of $\Lambda$.
@@
Suppose $x^{*}_{0}$ is in a (k2){(k-2)}-regular block $B$ of $\Lambda$. Since $B$ is $2$-connected, $y^{*}_{0}\notin V(B)$. If there exists a vertex $y_{1}^{*}\in N(V(B))\setminus(V(B)\cup\{y^{*}_{0}\})$, let $G^{\prime}=G(B,y^{*}_{0},y_{1}^{*})$. Recall that $y_{0}^{*},y_{1}^{*}\in V(G^{\prime})$. Since $y^{*}_{0}y_{1}^{*}$ is a cut edge in $G^{\prime}$, it is not part of an exceptional graph. Therefore, $\rho_{G^{\prime},h}(V(G^{\prime}))\leq-2$, and hence $\rho_{G,h}(V(G^{\prime}))\leq 2\lambda-2$. However, the multigraph $G[V(G^{\prime})]$ is disconnected. As $G[S^{*}_{0}]$ is an edge-block of $G$, $|E_{G}(V(G^{\prime})\cap S^{*}_{0},S^{*}_{0}\setminus V(G^{\prime}))|\geq 2$; therefore, by Lemma 4.5, $\rho_{G,h}(V(G^{\prime})\cap S^{*}_{0})\geq 2\lambda-3$. Hence, $\rho_{G,h}(V(G^{\prime})\setminus S^{*}_{0})\leq(2\lambda-2)-(2\lambda-3)<\lambda-1$, contradicting Observation 4.7.
Suppose now that $V(B)$ has no neighbor in $V(G)\setminus(V(B)\cup\{y^{*}_{0}\})$. Then every vertex $u\in V(B)-x_{0}^{*}$ satisfies h(u)=k2h(u)=k-2. Let $b=|V(B)|$, so
@@
If $k\geq 5$, then ${\Lambda}_{0}$ does not contain $K_{k-1}$ formed by (k1)(k-1)-vertices.
@@
Suppose there is a set $C=\{v_{1},v_{2},\ldots,v_{k-1}\}\subseteq S^{*}_{0}$ consisting of low (k1)(k-1)-vertices such that $B:=G[C]=K_{k-1}$. For $i\in[1,k-1]$, let $v^{\prime}_{i}$ be the unique neighbor of $v_{i}$ outside of $C$. Let $C^{\prime}=\{v_{1}^{\prime},v_{2}^{\prime},\ldots,v_{k-1}^{\prime}\}$, so that $v_{i}v_{i}^{\prime}\in E(G)$ for each $i\in[1,k-1]$. Since $G$ has no $K_{k}$, $|C^{\prime}|\geq 2$.
For distinct $v^{\prime}_{s},v^{\prime}_{t}\in C^{\prime}$, write $G_{s,t}=G(B,v_{s}^{\prime},v^{\prime}_{t})$. We will simultaneously consider two cases: If $\rho_{h}(G_{s,t})\leq-2$ for some distinct $v^{\prime}_{s},v^{\prime}_{t}\in C^{\prime}$, or if $G_{s,t}$ is exceptional for every pair of distinct $v^{\prime}_{s},v^{\prime}_{t}\in C^{\prime}$. Let $\varrho(C)=\min\{\rho_{h}(G_{s,t}):v^{\prime}_{s},v^{\prime}_{t}\in C^{\prime}\mbox{ and }v^{\prime}_{s}\neq v^{\prime}_{t}\}$.
For $j\in[1,k-1]$, define the function
[…]
For $j\in[1,k-1]$, let $\pi(j)$ be the minimum of $\rho_{G,h}(P)$ over $P\subseteq V(G)\setminus C$ with $|N(P)\cap C|\geq j$. Let $P_{j}\subseteq V(G)\setminus C$ be such that $|N(P_{j})\cap C|\geq j$ and $\rho_{G,h}(P_{j})=\pi(j)$.
We will prove that $\pi(k-1)\leq g(k-1)$. For this, let $j$ be the largest value for which $\rho_{G,h}(P_{j})\leq g(j)$. If j=k1j=k-1, then we are done, so assume j<k1j<k-1. Then $j\geq 2$ since
[…]
Assume there are distinct $i,i^{\prime}\in[1,k-1]$ such that $v^{\prime}_{i}\in P_{j}$ and $v^{\prime}_{i^{\prime}}\not\in P_{j}$. Since $V(G_{i,i^{\prime}})\cup P_{j}\supseteq P_{j}\cup\{v^{\prime}_{i^{\prime}}\}$, $V(G_{i,i^{\prime}})\cup P_{j}$ has at least $j+1$ neighbors in $C$. Thus, Lemma 2.2 implies that
(16) $\pi(j+1)\leq\rho_{G,h}(P_{j}\cup V(G_{i,i^{\prime}}))\leq g(j)+\rho_{G,h}(V(G_{i,i^{\prime}}))-\rho_{G,h}(P_{j}\cap V(G_{i,i^{\prime}})).$
@@
[…]
This contradicts Observation 4.6 in all cases except when $k=5$ and $\varrho(C)=5$. Since $1<\lambda-1$ when $k=5$, Observation 4.7 implies $C\cup P_{C^{\prime}}=V(G)$. In particular, $G[P_{C^{\prime}}]$ is $K_{5}^{-}$ and $G[C^{\prime}]$ is $K_{4}^{-}$ so $G_{s,t}$ is exceptional only for one pair $s,t$. Yet, this is impossible since we are in Case 2. ∎
@@
${\Lambda}_{0}$ does not contain a (k2){(k-2)}-regular block.